Solve the following problem and include your steps you use to solve this problem.
A roof truss is in the shape of a triangle with a height of x feet and a base of 2x + 1 feet. Write a polynomial that represents the area of the triangle. What is the area if x is 5 feet?

Answer by vidyamanohar(13)

Area is half the product of base and height.
So, Area=x(2x+1)/2=x^2+x/2
and when x=5, the area is 5(11)/2=55/2